### v1

@@ 7 Ranking of baseball players: An example by Efron and Morris
Figure 1: An alternative visualization of the Klein et al. (2020) 90% joint confidence region for overall ranking of baseball players with overlapping credible distributions for UB and HB models
This confidence set for the rank vector is depicted in Figure 1. Possible ranks for each player in this set are shown by a yellow line segment stretching from 1 to 18. According to this confidence set, any player can rank from 1 to 18, and any particular ranks can be held by any of these 18 players; clearly, it is totally uninformative. Due to multiple testing, the critical value $2.76$ (cf. equation (13) of Klein et al. (2020)) is too large to declare any comparison between two direct estimates to be significant. A joint $90\%$ credible distribution for the rank vector is a double-stochastic matrix, each row and each column of the matrix sums to one. This probability matrix is pictorially presented by overlaying on the figure of the confidence set solid red circles with area of a circle is proportional to the probability it is representing. The top matrix is based on the Cartesian credible set for $\theta$ under the unstructured Bayes model. From the figure, it is clear that many of the probabilities in this matrix are practically zeros implying that those combinations of players and ranks are improbable. Undoubtedly, the credible distribution is a lot more informative than the confidence set.
The observed rank $R_{i}$ of the $i$th player based on $y$ is shown by a blue square on the $R_{i}$th row and the $i$th column of the figure. If players $i_{1}$, $i_{2}$, $\cdots,i_{k}$ have the same $y$-values tied at ranks $r,r+1,\cdots,r+k-1$, we put $k$ blue squares on columns $i_{1}$, $i_{2}$, $\cdots,i_{k},$ of the $(r+k-1)$th row of this figure. [We followed the recommendation of Klein et al. (2020) to assign the highest possible rank to the tied players.] We used the remainder of the season performances of the players to compute their “true” rank vector $\xi$. The $i$th component, $\xi_{i}$, “true rank” for player $i$ is similarly depicted by a black circle on the $(\xi_{i},i)$ cell of the matrix.
The approach to the solution in the top panel of Figure 1 is the one which is the most similar to the Klein et al. (2020) solution. It is clear from the first row of this panel is that Player 1 has the largest probability of being the best hitter, and only the first five or six players have realistic chances to be the best hitter. Similarly, some of the players 12 through 18 are the most likely ones who had the lowest performance. These lists of players are much narrower than the whole set of $18$ players suggested by Klein et al. (2020) method. From the first column of this panel, on the other hand, we find that a realistic list of ranks for Player 1 is $\{14,15,16,17,18\}$. This list is much narrower than the entire set $\{1,2,\cdots,18\}$ suggested by the frequentist method. This figure also indicates that most likely ranks for player 17 are 1 to 7. The rank $17$ is most likely to be held by players 1 to 7. Clearly, the Bayesian solution is substantially more informative. In the bottom panel of Figure 1 we depicted the double stochastic matrix from the HPD credible set for $\theta$ using the HB model. Due to shrinkage of the joint posterior distribution towards a vector of common means, probability distributions of the rankings of the few uppermost and the lowermost players are not as sharp as those provided by the corresponding credible distributions from the UB model. Even though there is a weakness of the HB method for its reliance on shrinkage to a common mean in the absence of a useful covariate, the method provides better solutions than the UB method for some players whose early batting averages are at odds with their latter scores. Here, five players (4, 8, 13, 16, 17) showed a large or moderately large discordance between their early and late season performances. Two most discordant players are 4 and 17. Player 4 did not perform well in the remainder of the season which earned him a very low surrogate rank $3$ compared to his high $y$ rank $15$. From the top panel of Figure 1 we see that whereas the UB method assigns, perhaps erroneously, nearly 87% probability of Player 4 having a rank higher than 9, the bottom panel of Figure 1 reveals that the HB method correctly assigns less than 63% probability to this player having such a high rank. At the other end, Player 17 performed very well in the remainder of the season which earned him a very high surrogate rank $17$ (vs. a very low $y$-rank 2). Here, again, Figure 1 for the unweighted UB Cartesian method assigns, perhaps erroneously, nearly 93% probability of it having a rank lower than 10, the bottom panel of Figure 1 for the HB method correctly assigns less than 67% probability to it having such a low rank. Revised rankings of the discordant players afforded by the HB method are noteworthy and more reliable. We have demonstrated through simulation in Subsection 7.1 and the median income example in Section 9 that the HB method in the presence of useful covariates is substantially more effective. Also, for each player the variance of the possible ranks in the joint Klein et al. (2020) confidence region is $26.9$, the largest possible value, and the corresponding posterior variances from the credible distributions of ranks for these players vary between 5.36 and 16.36 for the UB model, and between 16.75 and 23.40 for the HB model. These numbers show a superiority of the credible distributions, created from the UB model or the HB model. We describe additional comparisons of the frequentist and the Bayesian solutions below. Recall that $\xi_{i}$ is the rank of $\omega_{i}$, which is a surrogate or a “gold standard” value of $\theta_{i}$. We use these gold standard values $\xi_{i}$ of the ranks $\check{r}_{i}$ to assess accuracy of various credible distributions of overall rankings. In particular, we compute for player $i=1,\cdots,m,$
@@
Table 1: Batting averages of 18 Major League players in 1970: $y_{i}$ is the average from first 45 at bats, $\omega_{i}$ the remainder of the season average, $R_{i}$ rank of $y_{i}$, $\xi_{i}$ rank of $\omega_{i}$, and deviation measures described in (14) for five overall ranking set estimators. For each of the two Bayesian models, we have two methods based on elliptical (subscript E) and Cartesian (subscript C) sets
1. Clemente .400 .346 18 18 5.90 5.90 1.18 1.18 8.50
2. F. Robinson .378 .298 17 15 4.39 4.39 1.87 1.87 6.17
3. F. Howard .356 .276 16 13 4.24 21.0 2.73 2.73 5.17
4. Johnstone .333 .222 15 3 8.11 8.11 11.40 11.39 6.83
5. Berry .311 .273 13.5 12 4.33 4.33 2.44 2.44 4.83
6. Spencer .311 .270 13.5 11 4.25 4.25 2.96 2.96 4.61
7. Kessinger .289 .263 12 7 4.96 4.96 4.92 4.92 4.83
8. L. Alvarado .267 .210 11 2 7.66 7.66 8.13 8.13 7.61
9. Santo .244 .269 9.5 10 4.36 4.34 3.09 3.09 4.50
10. Swoboda .244 .230 9.5 5 5.41 5.41 3.98 3.98 5.61
11. Unser .222 .264 6 8.5 4.26 4.26 3.09 3.08 4.56
12. Williams .222 .256 6 6 4.40 4.40 2.64 2.64 5.17
13. Scott .222 .303 6 16 7.64 7.64 9.36 9.36 6.83
14. Petrocelli .222 .264 6 8.5 4.25 4.25 3.33 3.33 4.56
15. E. Rodriguez .222 .226 6 4 5.38 5.38 3.46 3.46 6.17
16. Campenaris .200 .285 3 14 6.64 6.64 8.97 8.97 5.61
17. Munson .178 .316 2 17 9.54 9.54 13.57 13.56 7.61
18. Alvis .156 .200 1 1 5.46 5.46 1.43 1.43 8.50
Total abs. dev. 101.16 101.16 88.56 88.56 107.64
$10^{13}\times$ Vol. 51.38 1.32 93756 1757 85722
Average length 0.236 0.193 0.359 0.288 0.357
From Table 1 we find that the Klein et al. (2020) method has the highest sum of average absolute deviations, about 6.4% larger than for the HB method, and about 21.5% larger than for the UB method. This is counter-intuitive as the shrinkage estimators of $\theta_{i}$’s from the exchangeable HB model have lower total squared error than the same for the direct estimators $Y_{i}$’s, which are Bayes estimates of $\theta_{i}$ under the UB model. We have noted above that the total squared error of UB estimate of $\theta$ from the “true” vector $\omega$ was $0.07537$, and for the HB estimate it was $0.02368$. We present below a possible explanation for the domination of the UB method over the HB method when it comes to inference for overall ranking based on their credible distributions. In this example, estimating $\theta_{i}$ from the HB model shrinks the direct estimate; the shrinkage of $y_{i}$ towards a grand mean, measured roughly by the posterior mean of $E[D_{i}/(D_{i}+A)|y]$, is substantial (ranging between $0.75$ and $0.63$, and the shrinkage is more for the larger $D_{i}$’s that correspond to larger $y_{i}$’s. In our example, $D_{i}$’s decrease from $0.0053$ to $0.0029$, and the posterior mean of $A$ is $0.0024$ and the median of $A$ is $0.0017$.) These indicate that even in comparison with the small $D_{i}$’s, $A$ is small. In the absence of a covariate, a very small model error variance for an intercept only model essentially forces that $\theta_{i}$’s are all nearly equal. Accurate inference for the rank vector for such $\theta$ will naturally be challenging. Among the components of $\theta$, any differentiating information about $\theta_{i}$’s is given by its individual $y_{i}$’s only. Clearly, any shrinkage of an estimator of $\theta_{i}$ to a common value will reduce the estimator’s reliance on individual $y_{i}$’s, and thus the HB posterior distribution of $\theta_{i}$’s for ranking will be less informative than $y_{i}$, which is the center of the individual UB posterior distributions. Due to lack of a covariate, corresponding to the set of larger observations (the first 4 or 5 observations which are in descending order), the HB posterior distribution of any of these $\theta_{i}$’s will shrink downward (to some grand mean). This downward shrinking of the posterior distribution of $\theta_{i}$’s will lower the probability of them holding the higher ranks (18 through 15 or 14). In fact, from Table 2 the posterior expectation of $\check{r}_{1}$ in our example is $12.1$, which is quite a bit smaller than the largest rank $18$ (in this case) it is supposedly estimating. Similarly, for the next two smaller observations, the posterior expected ranks are again quite smaller than the target ranks 17 and 16. These discrepancies contribute to a bigger absolute deviations reported in Table 1. On the other hand, the distributions of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ for the UB model do not shrink downward much since the posterior of $\theta_{i}$’s are independent and each is concentrated around its $y_{i}$. The corresponding posterior means of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ are $16.8,16.2$ and $15.2$ (cf. Table 2), which are not much smaller than $18,17$ and $16$, respectively. These lesser discrepancies make smaller contributions to the expected absolute deviations reported Table 1. A similar shrinkage upward will be noticed for the set of smaller observations (observations 14 through 18). Corresponding expected absolute deviations of the rank from the surrogate true rank $\xi_{i}$ for observations 14 through 18 under the HB model again tend to be bigger than their counterparts from the UB model. We see bigger discrepancies between the HB ranks and the UB ranks for the first 4 and the last 4 observations. All of these mostly explain the smaller total posterior expected absolute rank deviations for the UB model in comparison with their counterparts for the HB model.
Table 2: Estimation of ranks of baseball players by different methods
Players P01 P02 P03 P04 P05 P06 P07 P08 P09
$y$ ranks 18 17 16 15 14 13 12 11 10
$\omega$ ranks 18 15 13 3 12 11 7 2 10
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50
HB-Cart 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
HB-Ellip 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
UB-Cart 16.82 16.20 15.23 14.40 12.98 12.80 11.51 10.12 8.40
UB-Ellip 16.82 16.20 15.23 14.39 12.98 12.79 11.51 10.12 8.40
Players P10 P11 P12 P13 P14 P15 P16 P17 P18 Average
$y$ ranks 9 8 7 6 5 4 3 2 1 9.5
$\omega$ ranks 5 8.5 6 16 8.5 4 14 17 1 9.5
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.5
HB-Cart 9.22 8.61 8.54 8.60 8.56 8.56 7.91 7.50 6.46 9.5
HB-Ellip 9.22 8.61 8.54 8.60 8.56 8.56 7.92 7.50 6.46 9.5
UB-Cart 8.41 6.71 6.69 6.64 6.43 6.79 5.03 3.43 2.43 9.5
UB-Ellip 8.41 6.72 6.69 6.64 6.43 6.79 5.03 3.44 2.43 9.5
Even though the UB method is poised to have smaller total expected absolute deviations than the HB method without a covariate, for some individual players the HB method is better suited to estimate ranks than the UB method. Continuing the above discussion further, since the posterior distribution of $\theta_{i}$ under the UB model is determined only by $y_{i}$, in case of a big discrepancy between $y_{i}$ and the unknown $\theta_{i}$ it is estimating (which will be reflected through a bigger difference between $y_{i}$ and $\omega_{i}$, the surrogate value of $\theta_{i}$), the posterior expected absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ will be large ($\check{r}_{i}$, the rank of $\theta_{i}$, and $\xi_{i}$, the rank of gold standard value $\omega_{i}$). Indeed this happens for five players 4, 8, 13, 16 and 17 with large discordance between a player’s early and late season performances. Among these five players, player 17 has the second smallest observation, who displayed large discrepancies between $y_{i}$ and $\omega_{i}$. Deviations $\epsilon_{i,UB}$ for these players are much larger than the average deviations of the other players, and these deviations also exceed the corresponding players’ deviations $\epsilon_{i,HB}$ based on the exchangeable HB model. If one or more player substantially over-perform or under-perform in the remainder of the reason, based on their later performances the true ranks of these players will change substantially from their ranks determined by $y_{i}$’s. The HB method by virtue of its borrowing from the other players is better suited to rectify the erroneous ranks of the players based on $y_{i}$’s. This is the likely reason that the HB ranks for the five players identified above perform better than the UB ranks that are based on individual $y_{i}$ alone; a downward shrinkage of the HB posterior distribution of the over-performing (in their first 45 at bats) players 4 and 8, and an upward shrinkage of the HB posterior distributions of the under-performing (in their first 45 at bats) players 13, 16 and 17. The HB posterior distributions of $\theta_{i}$ of these players, by virtue of their borrowing information from the other players, are substantially pulled toward the “true value”, measured by $\omega_{i}$. This is a useful property of the HB model. No such borrowing of information is possible through the UB model as the posterior distributions of the components $\theta_{i}$’s are independent and centered around $y_{i}$. Lastly, from this example where the total posterior expectation of absolute deviations of the true rank $\check{r}_{i}$ from the surrogate rank $\xi_{i}$ for the UB model is lower than the same for the HB model, we may not conclude that the UB model will always provide a more accurate solution. In fact, it is shown below by simulations that if a covariate effectively explains the $\theta_{i}$’s, the credible distributions of overall ranking based on the HB model are superior to the credible distributions derived from the UB model. Intuitively, this occurs since the posterior distribution of $\check{r}_{i}$ under the HB model will better concentrate near the true rank than the same posterior distribution under the UB model. We should note that even though in terms of total of expected absolute deviations, the frequentist Klein et al. (2020) method fares the worst, it is surprisingly robust for the five under-/over-performing players 4, 8, 13, 16 and 17. Finally, we included some size measures of the confidence set (for the KWW method) and credible sets for Bayesian methods, the sets are critical for the ranking problem. The last two rows of Table 1 provide volumes and approximate average lengths of the sets, respectively. In terms of volume, the HB sets are at least a thousand times better than the other sets. In terms of length, the HB sets are at least 50% better than the other sets. The best among all five sets is the HB elliptical set.
7.1 A simulation comparison of the set estimators of overall ranking

### v2

@@ 8 Ranking of the US states based on commuting times
Figure 1: An alternative visualization of the 90% joint confidence region for travel time ranking from Figure 1 of KWW with overlapping CDs from UB Cartesian sets (weighted)
We overlaid on Figure 1 the CDs based on a Cartesian credible set from the UB model. Both in terms of the model and the method, this table is most directly comparable with Figure 1 of Klein et al. (2020). Probabilities from CDs are shown by yellow circles, bigger circles for larger probabilities.
To clearly interpret the probabilities depicted in the figure, let us focus specifically on its $4$th row and $4$th column. The 4th column represents the state of NE has ranks $3-6$ with probabilities $0.21$, $0.20$, $0.55$ and $0.04$, respectively. We can contrast this with the red band on the 4th column in the figure which shows the rank set for NE suggested by the Klein et al. (2020) solution. The band shows that in addition to the four ranks $3-6$, NE can also rank 1, 2, 7 and 8.
On the other hand, the fourth row of the figure corresponding to rank 4 shows that the states that can plausibly hold this rank are AK, MT, NE and WY (with respective probabilities $0.04$, $0.23$, $0.55$ and $0.18$). Again, in contrast, the 4th row in Figure 1 of Klein et al. (2020) shows that states SD, ND, WY, NE, MT, AK, IA, KS and ID are all contenders of rank 4. Compared to our Bayesian CD solution, the frequentist solution includes five additional states. However, our CD for ID shows that it is virtually impossible for ID to be a contender of rank 4.
Again, if we consider ID, we find from Figure 1 that ID can hold rank 9 with probability virtually 1, and the rank 9 can only be held by ID. However, we observed it earlier that for the frequentist solution possible ranks for ID are $4-9$, and the states that may be ranked 9th are WY, MT, AK, IA, KS and ID. Similarly, while SD and ND, tied by direct estimates, both have the same frequentist rank set $1-6$, whereas in terms of Bayesian CDs, only probable ranks for these states are $1$ and $2$, and SD (ND) has $60\%\ (40\%)$ probability for rank $1$. Clearly, all these demonstrate that our Bayesian results are “less dispersed”, “more illuminating” and “more conclusive”.
On a related note, we observe that based on the facts that $y_{50}=y_{51}$ and they are the smallest among all $y$’s and $D_{51}<D_{50}$, it can be shown that for the HB model, $P(\check{r}_{51}>\check{r}_{50}|y)=P(\theta_{51}>\theta_{50}|y)<1/2$. Thus, we may take it as a justification that under the CDs $\{\check{r}_{51}=1\}$ is more likely than $\{\check{r}_{50}=1\}$; based on the HB CDs, we got these probabilities $\approx 0.6$ vs. $\approx 0.4$.
Table 1: Sizes CD-s-HB-C of (weighted) Cartesian HB CDs, KWW-c cardinality of KWW sets, entropies CD-e-HB-C of (weighted) Cartesian HB CDs, and KWW-e of KWW sets for ranks of nine states that have lower values of direct estimates of commuting times
States ID KS IA AK MT NE WY ND SD Average
CD-s-HB-C 1.00 2.23 1.93 3.84 3.46 2.31 4.49 1.95 1.95 2.57
KWW-c 6 7 7 9 9 8 9 6 6 7.44
CD-e-HB-C 0.01 0.74 0.88 1.39 1.33 1.12 1.42 0.69 0.68 0.92
KWW-e 1.79 1.95 1.95 2.20 2.20 2.08 2.20 1.79 1.79 1.99
We compare in Table 1 the KWW frequentist solution with the HB Cartesian CDs based on their “sizes” and “informativeness” using a subset of 9 states that have lower values of direct estimates of commuting times. First two rows measure the size and show that the CDs have smaller “size”. Last two lines measure the informativeness which show that the CDs are more “informative”.
For these 9 states, we also compute the credible distributions (i.e., the double stochastic matrix introduced in Section 5), shown in Table S.1 in Supplement, based on the elliptical region from the HB model. From the $4$th row of this matrix we find that the states AK, MT, NE and WY are the contenders of rank 4, and they may hold this rank with probabilities $0.11,0.21,0.50$ and $0.18$, respectively. The fourth column of that matrix conveys that NE can hold the ranks 3, 4, 5 and 6 with probabilities $0.32,0.50,0.15$ and $0.04$, respectively. Agreement between this set of probable values and those for the other cases such as the UB Cartesian show that credible regions of $\theta$ having different shapes and derived from different models provide reasonably similar solutions for overall ranking. From Table S.1 it is clear that with practical certainty that ID has rank 9, AK can have ranks $3-8$, and SD can have ranks 1 or 2. The corresponding KWW confidence sets, respectively, are $\{4-9\},\{1-9\}$ and $\{1-6\}$. This solution would assert that the ranks 4 for ID, 5 for AK and 6 for SD are plausible. However, this assertion would be strongly contradicted by the double scatter plot in Figure S.1. The plot makes it abundantly clear that commuting time for ID is higher than for AK, which in turn is higher than that for SD.
We have no covariate for this example and the HB estimate of the model error variance $A$, $13.2$ is very large; when put in perspective it is about $56$ times the largest sampling variance $D_{i}$ and $3572$ times the smallest one. With so high ratios of the model variance to sampling variances, the HB model is practically identical to the UB model. Similarity of the CDs from the UB and the HB models in this application are confirmed by the simulation study we presented in Table 4. In the last block of the table we set $A=1$ which is nearly $188$ times the largest sampling variance and $342$ times the smallest one. We find from the two lower-right-bottom blocks of this table that the average posterior mean absolute deviations reported for the UB and the HB columns show no practical differences among the four (HB and UB models for each of the Cartesian and elliptical credible sets) Bayesian solutions.
We also compare all five (one frequentist and four Bayesian) confidence and credible regions of $\theta$ in terms of their size measures. From Table 2 we find that all Bayesian credible sets except the UB-Cartesian one have smaller volumes than the KWW confidence set. While volumes of Cartesian credible sets are similar to the volume of the KWW set, the elliptical shape HPD credible sets from both models are enormously (by an order of $10^{-7}$) superior (this is due to cutting off low probability “corners” of the hyperrectangular sets to create an HPD set). Even in terms of average length, elliptical sets are better than the hyperrectangular sets. Although the KWW confidence intervals and UB Cartesian credible intervals are the same for $\theta$, when measuring the “sizes” of the KWW sets and the CDs for ranks, the average sizes are significantly smaller for the CDs (see the last line of Table 2).
Table 2: Commuting time: sizes of frequentist and Bayes set estimators of $\theta$ and $\check{r}$.
Frequentist Unstructured Bayes (weighted) Hierarchical Bayes (weighted)
Parameter Measure KWW Cartesian Elliptical Cartesian Elliptical
$\theta$ Volume 4.52 4.52 $1.66\times 10^{-7}$ 4.38 $1.56\times 10^{-7}$
$\theta$ Average length 1.16 1.16 0.83 1.16 0.83
$\check{r}$ Ave cardinality of rank sets 10.02 2.12 2.12 2.39 2.39
9 Ranking of the US states based on median incomes
Inference for median incomes of the US states for the income year 1989 has served as a benchmark example to evaluate effectiveness of various small area estimation methods. See for example, Ghosh et al. (1996) and Chung and Datta (2022). The Current Population Survey (CPS) provides annual estimates of median incomes for the states for various family sizes. These estimates serve as the direct estimates for the popular Fay-Herriot model. The application enables borrowing strength from other covariates by integrating administrative data and past census data. For this application two useful covariates are available. The first covariate, $x_{i1}$, is the $i$th state median income for 1979 from the 1980 Census. The second covariate, $x_{i2}$, is an adjustment to $x_{i1}$ for the $i$th state, $x_{i2}=(\text{PCI}_{i,1989}/\text{PCI}_{i,1979})x_{i1}$, $i=1,\ldots,m$, where $\text{PCI}_{i,1979}$ and $\text{PCI}_{i,1989}$ are the 1979 and 1989 per capita incomes of the $i$th state provided by the Bureau of Economic Analysis of the U.S. Department of Commerce. The 1989 census incomes for the states, which we denote by $\omega_{i}$, were medians based on a large number of households from each state. Due to large sample sizes many researchers (see, for example, Ghosh et al. (1996) and Chung and Datta (2022)) considered $\omega_{i}$’s as “gold standards” for true incomes, $\theta_{i}$, which we are trying to rank. We rank these surrogate values and treat their ranks, $\xi_{i}$, to be the surrogates for the true ranks $\check{r}_{i}$.
Figure 2: A visualization of the 90% joint confidence region for median income ranking by KWW with an overlapping credible distribution from the Cartesian HB method (weighted)
In Figure 2, we recreated Figure 1 of KWW that depicted the frequentist confidence region for ranking from equation (8). On this figure we overlaid the Cartesian credible distribution based on the weighted HB method. The KWW rank sets for all the states are depicted by appropriate yellow line segments that span a subset of 1 to 51. Probabilities from CDs are shown by red circles, bigger circles for larger probabilities. We use solid black circles on this figure to depict the “true” states ranks based on median incomes from the 1990 census. To the state CT the KWW solution assigns ranks 15 to 51, which is very wide. In contrast, the HB Cartesian CD assigns probabilities $0.03,0.13,0.33,$ and $0.51$, respectively to the ranks 48 to 51. For NJ the KWW ranks span from 22 to 51. The HB solution assigns probabilities $0.04,0.53,$ and $0.43$, respectively, to the ranks $49-51$. Clearly, the Bayesian CDs are very compact and informative compared to the frequentist solution.
Table 3: Median incomes: Comparison of set estimators of $\theta$ and $\check{r}$, derived by frequentist and Bayesian approaches (all weighted), by various measures
Parameter Measures $\epsilon_{i,\small{HB,C}}$ $\epsilon_{i,HB,E}$ $\epsilon_{i,UB,C}$ $\epsilon_{i,UB,E}$ $\epsilon_{i,KWW}$
$\check{r}$ Total abs. dev. 179.52 179.52 388.11 388.11 798.66
$\check{r}$ Ave entropy 2.27 2.27 2.62 2.62 3.86
$\check{r}$ Ave cardinality 10.27 10.27 15.57 15.57 47.94
$\theta$ Volume/$g$ $4.619\times 10^{7}$ $2.975$ $8.321\times 10^{14}$ $4.043\times 10^{7}$ $1.080\times 10^{15}$
[…]
In Table 3 we compare the frequentist and Bayesian solutions based on various measures. Line 1 compares the KWW solutions and the CDs using respective deviations from the surrogate values of ranks from the 1990 census based on equation (14). Line 2 compares these solutions in terms of their average entropies. Line 3 compares the solutions based on “sizes” of the rank solutions. Lines 4 and 5 list the scaled volumes and scaled ave lengths of the confidence/credible sets of $\theta$. Volumes of Bayesian sets (of $\theta$) are vastly smaller than those of the frequentist confidence sets. The HB elliptical set has the smallest volume. Average lengths of Bayesian sets are smaller than the average length of confidence set, again the HB elliptical set has the smallest average length.
Coming to line 1 of Table 3, we use the aforementioned surrogate values $\xi_{i}$ of the ranks $\check{r}_{i}$ to assess accuracy of various overall rankings. In particular, we compute for entity $i=1,\cdots,m,$
@@ 9 Ranking of the US states based on median incomes
In Table S.2 we presented for each state average absolute deviations (cf. equation (14)) of the KWW ranks from the surrogate ranks based on 1989 income data from the 1990 census. We also presented the posterior expected absolute deviations from (14) for the credible distributions (for the elliptic version only). Deviations for the KWW method exceed those for the HB method for all states, and $46$ KWW deviations except OK, GA, CO, NV, and DC exceed those for the UB method. We find that the KWW method has the highest sum of average absolute deviations (799), about 344% larger than for the HB method, and about 116% larger than for the UB method. Clearly, the best method is HB. The table shows that in terms of total of component expected absolute deviations of the rank solutions from the surrogate ranks (from equation (14), the CDs are vastly superior to the KWW solution. The HB CDs are better than the UB CDs but the shape of the credible sets for $\theta$ appears not to play any role. Similar conclusions emerge from the other two measures in Table 3 for evaluating solutions for $\check{r}$.
For the median income application we further explore comparison of credible distributions generated by various Bayesian models (UB and HB, with and without covariates) and methods (weighted and unweighted) among themselves and with the frequentist method of KWW in terms of various deviations of equation (14). We plotted several of these posterior expected absolute deviations against the KWW average absolute deviations in Figure 3.
The left panel in the first row of Figure 3 is the plot using posterior expected absolute deviations from credible distributions from weighted method based on the HB model with covariates on the $y$-axis and the KWW deviations on the $x$-axis. We use green circles for the elliptical method and black stars for the rectangular method. This panel shows that all the points cluster below the $y=x$ line and closer to the $x$-axis, indicating that the credible distributions are vastly superior to the KWW method. It also shows that for the weighted method there are practically no differences between the deviations from credible distributions based on the rectangular or the elliptical credible sets for the mean vector $\theta$.
The right panel in the first row of that figure differs from the left panel only by use of no covariates in the HB model. We see that several points are above the $y=x$ line, indicating that the KWW solution for those states are better than the HB solution in the absence of covariates. Also, although many of these points that are below the $y=x$ line, a large number of points have $y$-values bigger than 5, in contrast with the left panel where only about $15\%$ of the points are that big, indicating that the credible distributions without covariates are not as good as those created by using good covariates. These two panels convincingly demonstrate the importance of effective covariates for the ranking problem.
The left panel on the second row mimics the left panel on the first row except that now the credible distributions are unweighted. The HB model employs covariates. As we have noticed in the first row, in this case also we see that the Bayesian solutions (both elliptical and rectangular) are vastly better than the KWW solution. However, while most of the points on the panel in the second row have $y$-values bigger than 5, most are below that in the panel on the first row. These two panels show the superiority of the weighted credible distributions. In the right panel on this row we plotted the deviations based on weighted, elliptical credible distributions for the UB model and the HB model without covariates against the KWW deviations. While overall the credible distributions show superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates. This plot also shows that in the absence of a covariate the HB method loses its advantage over the UB method, and the two credible distributions performed rather similarly.
In the last row of Figure 3, we compare various CDs (weighted vs. unweighted, elliptical vs. Cartesian) from the UB model against the confidence set provided by the KWW method in terms of posterior expected absolute deviations and average absolute deviations resulting from the KWW method (computed from (14)). While again the UB credible distributions show overall superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates.
Figure 3: Various plots of expected or average absolute deviations for median incomes
Refer to caption Refer to caption
With Covariates Covariates not used
Refer to caption Refer to caption
With Covariates No covariates, elliptical
Refer to caption Refer to caption
Covariates cannot be used Covariates cannot be used
10 Ranking of baseball players: An example by Efron and Morris (1975)
Efron and Morris (1975) considered batting averages of 18 major league baseball batters from their first 45 at bats in the 1970 baseball season to predict their performances in the remainder of that season. Let $\theta_{i}$ be the unknown true batting average of player $i$ during that season. A reasonable representative value for $\theta_{i}$ was the player’s average ($\omega_{i}$) in the remainder of the season. The sample proportion of hits $Y_{i}$ from first 45 at bats is approximately normal with mean $\theta_{i}$ and variance $D_{i}$. In our analysis, we use the estimated variance, given by $D_{i}=Y_{i}(1-Y_{i})/45.$ Treating $\omega_{i}$ as the “gold standard” value of $\theta_{i}$, the total empirical squared error (TESE), $\sum_{i=1}^{18}(y_{i}-\omega_{i})^{2}$, for the direct estimates $y_{i}$’s, is $0.07537$, whereas $0.02368$ for the HB, and $0.02356$ for the empirical best linear unbiased predictor (EBLUP). The TESE for the direct estimates is about $3.2$ times of that of HB or the EBLUP estimates.
Efron and Morris (1975) used the arc-sin transformed parameters $\psi_{i}=arcsin(2\theta_{i}-1)$, for the transformed data $\hat{\psi}_{i}=arcsin(2Y_{i}-1)$ in their EB analysis. This transformation resulted in equal approximate sampling variances, $1/45$, for all $\hat{\psi}_{i}$’s, $i=1,\cdots,18$. They applied Stein’s shrinkage estimation method based on an exchangeable model to calculate ${\hat{\psi}}_{J-S,i}$, the James-Stein estimate of $\psi_{i}$. The TESE of prediction for the direct estimates $\hat{\psi}_{i}$ from $arcsin(2\omega_{i}-1)$ is 0.3902, which is about 3.5 times of 0.1113, the TESE for the James-Stein estimates ${\hat{\psi}}_{J-S,i}$. This application showed a substantial efficiency of Stein’s method. The efficiency for HB estimators of $\psi_{i}$ based on the same model is about $3.4$, which is comparable. Note that these numbers compare well with $3.2$, we obtained above for EBLUP/HB predictors of $\theta_{i}$’s.
Figure 4: A visualization of the KWW 90% joint confidence region for overall ranking of baseball players with overlapping CDs for UB and HB models (weighted, elliptical method)
This confidence set for the rank vector is depicted in Figure 4. Possible ranks for each player in this set are shown by a yellow line segment stretching from 1 to 18. According to this confidence set, any player can rank from 1 to 18, and any particular rank can be held by any of these 18 players; clearly, results are totally uninformative. Due to multiple testing, the critical value $2.76$ from equation (13) of Klein et al. (2020) is too large to declare any comparison between two direct estimates to be significant. A joint $90\%$ credible distributions for the rank vector is a double-stochastic (DS) matrix. This DS matrix is pictorially presented by overlaying on the figure of the confidence set solid red circles with area proportional to the probability it is representing. The top matrix is based on the elliptic set for $\theta$ under the UB model and the bottom matrix is a similar version for the intercept only HB model (both weighted).
From the top panel (UB) of Figure 4, it is clear that many probabilities in this matrix are practically zeros implying that those combinations of players and ranks are improbable. Undoubtedly, the resulting credible distributions are lot more informative than the frequentist confidence set. The bottom panel of this figure, corresponding to the HB intercept only, again shows that many red circles are still tiny, meaning that those combinations are less likely. However, in comparison with the top panel, these circles, or the probabilities, are bigger. This means that the CDs in the intercept case are less concentrated. This happens due to shift of the posterior distribution of each true mean $\theta_{i}$ to a common grand value (estimate of the intercept). Due to space limitation, a full discussion of this example is moved to Supplemental Section S.4.
In this example using the surrogate ranks $\xi_{i}$’s for the players, we computed the posterior expected absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ using equation (14) for both the UB and the HB models. These values for all players are displayed in Table S.3. Thirteen of the eighteen deviations for the UB model-based solution are smaller than those for the HB model. Total of these deviations for the UB model are also smaller than its HB counterpart. The HB model performs worse. In contrast, the shrinkage estimates of $\theta$ (HB estimates based on the intercept model) produced a far less TESE, $0.02368$, compared with $0.07537$, the TESE of the UB model-based estimates of $\theta_{i}$’s. In the absence of any covariate, the HB model-based estimation of $\theta$ and ranks $\check{r}_{i}$’s go in the opposite directions. Similar results were obtained by Goldstein and Spiegelhalter (1996) (see their Tables 6b and 7b). But if good covariates are available that effectively explain the variability of $\theta$ (implied by small model error variance $A$ in the Fay-Herriot model), the credible distributions from HB models perform better than the CDs from an UB model. Simulation results that we report in Subsection 10.1 illustrate this point. A careful reading of the first two rows in each of the first three blocks for $A=0.001,0.005$ and $0.10$ in Table 4 clearly show this point. These results also shed light on why the CDs from the HB model in the median income example do significantly better than those from the UB model (cf. Figure 3 in Section 9).
10.1 A simulation comparing set estimators of overall ranking